# Disentangling Rich Dynamics from Feature Learning: A Framework for Independent Measurements

### v2

@@ C.2 Inductive bias toward large features for minimum norm solution
For overparameterized linear models, the inductive bias determines the returned function as many functions can express the training set. Here, we brief the linear models’ inductive bias toward larger (significant) features or eigenfunctions [16, 17, 18, 19, 20, 21, 54].
The inductive bias is quantified through learnability $L_{k}$ [54] or the expectation ratio between the learned coefficient and true coefficient for $e_{k}$:
$L_{k}:=\mathbf{E}_{S\sim q^{n}}\left[\frac{\braket{\hat{f}}{e_{k}}}{\braket{f^{*}}{e_{k}}}\right]=\frac{\rho_{k}}{\rho_{k}+\kappa},\quad\quad\textrm{where }\sum_{k=1}^{p}\frac{\rho_{k}}{\rho_{k}+\kappa}=n.$ (15)
@@
We can understand the inductive bias from the (vanishing) regularization on $L^{2}$ norm ($\lambda\|w\|_{F}^{2}$ in Eq. 11) and parameterization in Eq. 14. If different eigenfunctions can equally express the training set, eigenfunction with larger eigenvalue often requires smaller coefficient $\theta_{k}$ to express the samples. Because $\|w\|_{F}^{2}=\|\theta\|_{F}^{2}$, expressing the data with larger (eigenvalue) eigenfunctions minimizes the norm, creating an inductive bias toward larger features.
Instead of the derivation found in the references [16, 17, 18, 19, 20, 21] — which requires random matrix theory or replica trick — we show an example of the inductive bias in Fig. 9.
Figure 9: Different inductive biases by the eigenvalues (intensity). Two $10$-parameter linear models are trained on 4 datapoints (overparmaterized) with gradient flow. Both linear models use $\sin$ basis functions such that $f(x)=\sum_{k}^{10}w_{k}k^{\alpha}\sqrt{2}\sin(2\pi kx)$ — spanning an identical function space — but differ in the eigenvalues with $\alpha=-4$ (blue) and $\alpha=4$ (orange), leading to different inductive biases. (a): The blue model has a greater intensity for lower-frequency $\sin$ functions, while the orange model shows the opposite. (b): The learned functions show that the blue model used lower-frequency functions while the orange used higher-frequency functions to express the training samples.
Fig. 9 shows that two linear models with same hypothesis space but different eigenvalues learn dramatically different functions. The blue model has large eigenvalues for low-frequency functions and fits the datapoints using low-frequency functions. The orange model with opposite eigenvalue distribution fits the training set with high-frequency functions — a clear example of an inductive bias toward large (intensity) features.
@@
$\mathcal{L}_{G}=\frac{1}{1-\frac{1}{n}\sum_{j}^{p}L_{j}^{2}}\left(\sum_{k}^{p}(1-L_{k})^{2}\braket{e_{k}}{f^{*}}^{*2}\right).\$ (16)
The equation formalizes the task-model alignment: the alignment between large features (large $\rho_{k}$ thus large $L_{k}$) and high quality (large $\braket{e_{k}}{f^{*}}$) leads to better generalization.
@@ C.3 Dynamical inductive bias toward large features
where in the second line, we used Eq. 11 on infinite datapoints so $\mathcal{L}_{emp}=\mathcal{L}$. In the last line, we used the definition of integral operator (Eq. 1).
Using the eigenvalues $\rho_{k}$ and eigenfunctions $e_{k}$ of the kernel integral operator $\mathcal{T}$, we obtain the dynamics for each $\braket{e_{k}}{f}$ by inner producting $e_{k}$ on both sides of Eq. 17,
$\frac{d\braket{e_{k}}{f}}{dt}=\mathbf{E}_{x^{\prime}}\left[e_{k}(x^{\prime})\frac{df}{dt}(x^{\prime})\right]$ $=-\mathbf{E}_{x^{\prime}}\left[e_{k}(x^{\prime})\mathcal{T}\left[f-f^{*}\right](x^{\prime})\right]$ (22)
$=-\braket{e_{k}}{\mathcal{T}[f-f^{*}]}$ (23)
$=-\rho_{k}(\braket{e_{k}}{f}-\braket{e_{k}}{f^{*}}),$ (24)
where we use the definition of eigenfunction in the last line (Eq. 2). Expanding $f$ in the eigenfunction basis and plugging in Eq. 22, the function $f$ at time $t$ is a sum of $p$ independent modes that saturate faster for larger eigenvalues:
$f(x)|_{t}=\sum_{k=1}^{p}\braket{e_{k}}{f^{*}}(1-e^{-\rho_{k}t})e_{k}(x),$ (25)
where we assumed $f$ is a zero function at initialization. Eq. 25 shows that gradient flow decouples the dynamics of linear models into $p$ modes, where each mode corresponds to the evolution of $\braket{e_{k}}{f}$ having a saturation speed governed by $\rho_{k}$ — thus the dynamical inductive bias toward larger eigenfunction.
@@ C.4 Application to our method
Here, we detail the intuition of our visualization methods. As described in Bengio et al.[1], feature learning (of representation) is providing better features for a simpler learner. The analogy allows us to describe a neural network as (1) feature map $\Phi$ providing better features for (2) the last layer — the simple learner (Fig. 10). Yet, better features for the last layer is ill-defined.
The last layer interacts with rest of the network only through the features. While the exact dynamics differs from linear models, the last layer maintains the dynamical inductive bias (Section C.3) toward larger features. Furthermore, the quality $(\braket{e_{k}}{f^{*}})$, utilization $(\braket{e_{k}}{\hat{f}})$, and intensity $\rho_{k}$ formalisms readily extend to the last layer features, motivating our visualization method.
Figure 10: A neural network decomposed into a feature map and a linear last layer. (left):An abstract diagram depicting a DNN architecture as a combination of the feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$ from the input space $\mathcal{X}$ to the $p$ post activations of the penultimate layer, and final linear classifier for $C$-way classifcation. Most neural networks share this abstract structure and mainly vary in how they create their feature maps. (right): Illustration of the visualization methods in function space.
Figure 11: Toy model demonstrating feature learning. A 4-layer MLP with width 1000 and scalar input and output is trained to learn the Heaviside step function $f^{*}$ over the domain $[-1,1]$. (a): The first three eigenfunctions (dashed) are shown at initialization, resembling orthogonal polynomials. (b,c): During training, features evolve such that a single feature (red) fits both the target and learned functions (grey and green). (d,e,f): Our visualization metrics defined in Eq. 5 at different epochs. (d) and (e) show that fewer features can express the target/learned function, and (f) shows that only a few features are significant after training.
In the rich regime, features evolve during training, and Fig. 11(a-c) demonstrates an example of change in features for 4-layer MLP trained to fit the Heaviside step function. The eigenfunctions are difficult to visualize in practice as the input space is high dimensional. Fig. 11(d-f) shows our visualization, capturing the evolution of features by measuring properties natural to linear regression.
@@ Appendix D The relationship between the MP-operator $\mathbf{\mathcal{T}_{MP}}$ and neural collapse
Neural collapse [6] (NC) refers to a state of a DNN when the feature vectors of the training set and last layer weights form a symmetric and clustered structure at the terminal phase of training (TPT) or when trained past the point where the training error vanishes. The structure has been studied mainly in the Unconstrained Feature Model (UFM) [5] and has sparked theoretical studies. See Kothapalli et al. [35] for a review.
@@
    $\Sigma_{W}\Sigma_{b}^{\dagger}\rightarrow 0,$ (26)
@@ D.1 Proofs of proposition 1 and 2
$\mu_{i}=\mathbf{E}_{x\in A_{i}}[\Phi(x)]=\frac{C}{n}\sum_{x\in A_{i}}\Phi(x)=C\braket{\hat{f}_{i}}{\Phi}.$ (31)
Note that $\braket{\hat{f}_{i}}{\Phi}\in\mathbb{R}^{p}$ is a vector with $\braket{\hat{f}_{i}}{\Phi}=[\braket{\hat{f}_{i}}{\Phi_{1}},\ldots,\braket{\hat{f}_{i}}{\Phi_{p}}]$.
@@
$\bar{\mu}=\frac{1}{C}\sum_{i}^{C}\mu_{i}=\braket{\Phi}{\sum_{i=1}^{C}\hat{f}_{i}}=\braket{\mathbf{1}}{\Phi}.$ (32)
@@
$\braket{\hat{f}_{i}}{\hat{f}_{j}}=\frac{1}{C}\delta_{ij},\quad\quad\sum_{i=1}^{C}\hat{f}_{i}=\mathbf{1},\quad\quad\braket{\mathbf{1}}{\hat{f}_{j}}=\frac{1}{C}.$ (36)
@@
$\braket{\mathbf{1}}{C\hat{f}_{i}-\mathbf{1}}=0.$ (37)
@@
$\Phi(x^{\prime})^{T}(\mu_{i}-\bar{\mu})$ $=\Phi(x^{\prime})^{T}\left(\braket{\Phi}{C\hat{f}_{i}}-\braket{\Phi}{\mathbf{1}}\right)$ (38)
@@
where we used the definition of the means (Eqs. 31 and 32) in the first line and the definition of $\mathcal{T}$ (Eq. 1) in the last line. As $\mathcal{T}$ is $\mathcal{T}_{MP}$, we obtain
@@
where we used Eq. 37 and the definition of $\mathcal{T}_{MP}$ (Definition 1). ∎
@@
In the setup of Lemma 1, we have
@@
From the definition of class mean (Eq. 31) and Lemma 1,
@@
We will show that $Tr(\Sigma_{b}\Sigma_{W})=0$, which automatically proves the proposition as covariance matrices are positive semi-definite. We have
$Tr(\Sigma_{b}\Sigma_{W})$ $=Tr\left(\sum_{j}(\mu_{j}-\bar{\mu})(\mu_{j}-\bar{\mu})^{T}\sum_{i}\mathbf{E}_{x\in A_{i}}(\Phi(x)-\mu_{i})(\Phi(x)-\mu_{i})^{T}\right)$ (46)
@@
where we used linearity of trace, sum, expectation in the first line and Lemma 1 in the second line.
Expanding Eq. 49,
$Tr(\Sigma_{b}\Sigma_{W})$ $=a_{2}^{2}C^{2}\sum_{i,j}\mathbf{E}_{x\in A_{i}}\left[(1-2\delta_{ij})\hat{f}_{j}(x)+\delta_{ij}\right]$ (50)
@@
From the definition of global mean (Eq. 32) and Lemma 1,
@@
D.2 The first eigenfunction, the constant function, the analog of global mean vector in simplext ETF
Here, we discuss the constant function, which, by definition of $\mathcal{T}_{MP}$ (Definition 1), is the largest eigenfunction if $\mathcal{T}$ is $\mathcal{T}_{MP}$. In all experiments except Fig. 11, the largest eigenfunction is the constant function throughout training, with $\braket{\mathbf{1}}{e_{1}}>0.95$. For our visualization, we ignore the first eigenvalue from analysis as the constant function is removed in CKA (Eq. 7).
In the NC2 condition, the $k$-simplex ETF structure is a set of $k$ orthogonal vectors projected on $k-1$ dimensional space along the compliment of the global mean vector. For example, projecting $[0,0,1]$, $[0,1,0]$, and $[1,0,0]$ onto the orthogonal complement of $[\frac{1}{3},\frac{1}{3},\frac{1}{3}]$ (the global mean vector) gives three vertices $[-\frac{1}{3},-\frac{1}{3},\frac{2}{3}]$, $[-\frac{1}{3},\frac{2}{3},-\frac{1}{3}]$, $[\frac{2}{3},-\frac{1}{3},-\frac{2}{3}]$ which forms an (ordinary plane) equilateral triangle.
Our operator $\mathcal{T}_{MP}$ (Definition 1) has an analogous structure, where it is a (scaled) projection operator ($a_{2}P_{\hat{\mathcal{H}}}$) except along the constant function. The $CKA$ (Eq. 7) also measures the alignment except along the mean direction (for matrices) or the constant direction (for operators). This allows neural networks to have any arbitrary eigenvalue for the constant function, yet perfectly align with $\mathcal{D}_{LR}=0$.
When $\mathcal{T}$ is $\mathcal{T}_{MP}$, the definition of $\mathcal{T}_{MP}$ (Definition 1) trivially shows that the largest eigenfunction is the constant function. A peculiar observation is that the first eigenfunction of $\mathcal{T}$ remains the constant function ($\braket{\mathbf{1}}{e_{1}}>0.95$) throughout training in all our experiments (except Fig. 11). This pattern also appears in hierarchical datasets [55], suggesting that natural data may consistently prioritize the constant function as the dominant mode — warranting further investigation.

### v3

@@ C.2 Inductive bias toward large features for minimum norm solution
For overparameterized linear models, the inductive bias determines the returned function as many functions can express the training set. Here, we brief the linear models’ inductive bias toward larger (significant) features or eigenfunctions (Bordelon et al., 2020; Jacot et al., 2020; Spigler et al., 2020; Canatar et al., 2021; Cui et al., 2021; El Harzli et al., 2024; Simon et al., 2023).
The inductive bias is quantified through learnability $L_{k}$ (Simon et al., 2023) or the expectation ratio between the learned coefficient and true coefficient for $e_{k}$:
$L_{k}:=\mathbf{E}_{S\sim q^{n}}\left[\frac{\braket{\hat{f}|e_{k}}}{\braket{f^{*}|e_{k}}}\right]=\frac{\rho_{k}}{\rho_{k}+\kappa},\quad\quad\textrm{where }\sum_{k=1}^{p}\frac{\rho_{k}}{\rho_{k}+\kappa}=n.$ (15)
@@
We can understand the inductive bias from the (vanishing) regularization on $L^{2}$ norm ($\lambda\|w\|_{F}^{2}$ in Eq.˜11) and parameterization in Eq.˜14. If different eigenfunctions can equally express the training set, eigenfunction with larger eigenvalue often requires smaller coefficient $\theta_{k}$ to express the samples. Because $\|w\|_{F}^{2}=\|\theta\|_{F}^{2}$, expressing the data with larger (eigenvalue) eigenfunctions minimizes the norm, creating an inductive bias toward larger features.
Instead of the derivation found in the references (Bordelon et al., 2020; Jacot et al., 2020; Spigler et al., 2020; Canatar et al., 2021; Cui et al., 2021; El Harzli et al., 2024) — which requires random matrix theory or replica trick — we show an example of the inductive bias in Fig.˜7.
Figure 7: Different inductive biases by the eigenvalues (intensity). Two $10$-parameter linear models are trained on 4 datapoints (overparameterized) with gradient flow. Both linear models use $\sin$ basis functions such that $f(x)=\sum_{k}^{10}w_{k}k^{\alpha}\sqrt{2}\sin(2\pi kx)$ — spanning an identical function space — but differ in the eigenvalues with $\alpha=-4$ (blue) and $\alpha=4$ (orange), leading to different inductive biases. (a): The blue model has a greater intensity for lower-frequency $\sin$ functions, while the orange model shows the opposite. (b): The learned functions show that the blue model used lower-frequency functions while the orange used higher-frequency functions to express the training samples.
Fig.˜7 shows that two linear models with same hypothesis space but different eigenvalues learn dramatically different functions. The blue model has large eigenvalues for low-frequency functions and fits the datapoints using low-frequency functions. The orange model with opposite eigenvalue distribution fits the training set with high-frequency functions — a clear example of an inductive bias toward large (intensity) features.
@@
$\mathcal{L}_{G}=\frac{1}{1-\frac{1}{n}\sum_{j}^{p}L_{j}^{2}}\left(\sum_{k}^{p}(1-L_{k})^{2}\braket{e_{k}|f^{*}}^{*2}\right).\\$ (16)
The equation formalizes the task-model alignment: the alignment between large features (large $\rho_{k}$ thus large $L_{k}$) and high quality (large $\braket{e_{k}|f^{*}}$) leads to better generalization.
@@ C.3 Dynamical inductive bias toward large features
where in the second line, we used Eq.˜11 on infinite datapoints so $\mathcal{L}_{emp}=\mathcal{L}$. In the last line, we used the definition of integral operator (Eq.˜1).
Using the eigenvalues $\rho_{k}$ and eigenfunctions $e_{k}$ of the kernel integral operator $\mathcal{T}$, we obtain the dynamics for each $\braket{e_{k}|f}$ by taking inner products with $e_{k}$ on both sides of Eq.˜17,
$\frac{d\braket{e_{k}|f}}{dt}=\mathbf{E}_{x^{\prime}}\left[e_{k}(x^{\prime})\frac{df}{dt}(x^{\prime})\right]$ $=-\mathbf{E}_{x^{\prime}}\left[e_{k}(x^{\prime})\mathcal{T}\left[f-f^{*}\right](x^{\prime})\right]$ (22)
$=-\braket{e_{k}|\mathcal{T}[f-f^{*}]}$ (23)
$=-\rho_{k}(\braket{e_{k}|f}-\braket{e_{k}|f^{*}}),$ (24)
where we use the definition of eigenfunction in the last line (Eq.˜2). Expanding $f$ in the eigenfunction basis and plugging in Eq.˜22, the function $f$ at time $t$ is a sum of $p$ independent modes that saturate faster for larger eigenvalues:
$f(x)|_{t}=\sum_{k=1}^{p}\braket{e_{k}|f^{*}}(1-e^{-\rho_{k}t})e_{k}(x),$ (25)
where we assumed $f$ is a zero function at initialization. Eq.˜25 shows that gradient flow decouples the dynamics of linear models into $p$ modes, where each mode corresponds to the evolution of $\braket{e_{k}|f}$ having a saturation speed governed by $\rho_{k}$ — thus the dynamical inductive bias toward larger eigenfunction.
@@ C.4 Application to our method
Here, we detail the intuition of our visualization methods. As described in Bengio et al. (2013), feature learning (of representation) is providing better features for a simpler learner. The analogy allows us to describe a neural network as (1) feature map $\Phi$ providing better features for (2) the last layer — the simple learner (Fig.˜8). Yet, better features for the last layer is ill-defined.
[…]
Figure 8: A neural network decomposed into a feature map and a linear last layer. (left):An abstract diagram depicting a DNN architecture as a combination of the feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$ from the input space $\mathcal{X}$ to the $p$ post activations of the penultimate layer, and final linear classifier for $C$-way classification. Most neural networks share this abstract structure and mainly vary in how they create their feature maps. (right): Illustration of the visualization methods in function space.
Figure 9: Toy model demonstrating feature learning. A 4-layer MLP with width 1000 and scalar input and output is trained to learn the Heaviside step function $f^{*}$ over the domain [1,1][-1,1]. (a): The first three eigenfunctions (dashed) are shown at initialization, resembling orthogonal polynomials. (b,c): During training, features evolve such that a single feature (red) fits both the target and learned functions (grey and green). (d,e,f): Our visualization metrics defined in Eq.˜5 at different epochs. (d) and (e) show that fewer features can express the target/learned function, and (f) shows that only a few features are significant after training.
In the rich regime, features evolve during training, and Fig.˜9(a-c) demonstrates an example of change in features for 4-layer MLP trained to fit the Heaviside step function. The eigenfunctions are difficult to visualize in practice as the input space is high dimensional. Fig.˜9(d-f) shows our visualization, capturing the evolution of features by measuring properties natural to linear regression.
@@ Appendix D The relationship between the MP-operator $\mathbf{\mathcal{T}_{MP}}$ and neural collapse
Neural collapse (Papyan et al., 2020) (NC) refers to a state of a DNN when the feature vectors of the training set and last layer weights form a symmetric and clustered structure at the terminal phase of training (TPT) or when trained past the point where the training error vanishes. The structure has been studied mainly in the Unconstrained Feature Model (UFM) (Mixon et al., 2020) and has sparked theoretical studies. See  Kothapalli et al. (2022) for a review.
@@
    $\Sigma_{b}^{\dagger}\Sigma_{W}\rightarrow 0,$ (26)
@@ D.1 Proofs of proposition 1 and 2
$\mu_{i}=\mathbf{E}_{x\in A_{i}}[\Phi(x)]=\frac{C}{n}\sum_{x\in A_{i}}\Phi(x)=C\braket{\hat{f}_{i}|\Phi}.$ (31)
Note that $\braket{\hat{f}_{i}|\Phi}\in\mathbb{R}^{p}$ is a vector with $\braket{\hat{f}_{i}|\Phi}=[\braket{\hat{f}_{i}|\Phi_{1}},\ldots,\braket{\hat{f}_{i}|\Phi_{p}}]$.
@@
$\bar{\mu}=\frac{1}{C}\sum_{i}^{C}\mu_{i}=\braket{\Phi|\sum_{i=1}^{C}\hat{f}_{i}}=\braket{\mathbf{1}|\Phi}.$ (32)
@@
$\braket{\hat{f}_{i}|\hat{f}_{j}}=\frac{1}{C}\delta_{ij},\quad\quad\sum_{i=1}^{C}\hat{f}_{i}=\mathbf{1},\quad\quad\braket{\mathbf{1}|\hat{f}_{j}}=\frac{1}{C}.$ (36)
@@
$\braket{\mathbf{1}|C\hat{f}_{i}-\mathbf{1}}=0.$ (37)
@@
$\Phi(x^{\prime})^{T}(\mu_{i}-\bar{\mu})$ $=\Phi(x^{\prime})^{T}\left(\braket{\Phi|C\hat{f}_{i}}-\braket{\Phi|\mathbf{1}}\right)$ (38)
@@
where we used the definition of the means (Eqs.˜31 and 32) in the first line and the definition of $\mathcal{T}$ (Eq.˜1) in the last line. As $\mathcal{T}$ is $\mathcal{T}_{MP}$, we obtain
@@
where we used Eq.˜37 and the definition of $\mathcal{T}_{MP}$ (Definition˜1). ∎
@@
In the setup of Lemma˜1, we have
@@
From the definition of class mean (Eq.˜31) and Lemma˜1,
@@
We will show that $\mathrm{Tr}(\Sigma_{b}\Sigma_{W})=0$, which automatically proves the proposition as covariance matrices are positive semi-definite. We have
$\mathrm{Tr}(\Sigma_{b}\Sigma_{W})$ $=\mathrm{Tr}\left(\sum_{j}(\mu_{j}-\bar{\mu})(\mu_{j}-\bar{\mu})^{T}\sum_{i}\mathbf{E}_{x\in A_{i}}(\Phi(x)-\mu_{i})(\Phi(x)-\mu_{i})^{T}\right)$ (46)
@@
where we used linearity of trace, sum, expectation in the first line and Lemma˜1 in the second line.
Expanding Eq.˜49,
$\mathrm{Tr}(\Sigma_{b}\Sigma_{W})$ $=a_{2}^{2}C^{2}\sum_{i,j}\mathbf{E}_{x\in A_{i}}\left[(1-2\delta_{ij})\hat{f}_{j}(x)+\delta_{ij}\right]$ (50)
@@
From the definition of global mean (Eq.˜32) and Lemma˜1,
@@
D.2 The first eigenfunction, the constant function, the analog of global mean vector in simplex ETF
Here, we discuss the constant function, which, by definition of $\mathcal{T}_{MP}$ (Definition˜1), is the largest eigenfunction if $\mathcal{T}$ is $\mathcal{T}_{MP}$. In all experiments except Fig.˜9, the largest eigenfunction is the constant function throughout training, with $\braket{\mathbf{1}|e_{1}}>0.95$. For our visualization, we ignore the first eigenvalue from analysis as the constant function is removed in CKA (Eq.˜7).
In the NC2 condition, the $k$-simplex ETF structure is a set of $k$ orthogonal vectors projected on k1k-1 dimensional space along the complement of the global mean vector. For example, projecting $[0,0,1]$, $[0,1,0]$, and $[1,0,0]$ onto the orthogonal complement of $\frac{1}{3},\frac{1}{3},\frac{1}{3}]$ (the global mean vector) gives three vertices $\frac{1}{3},-\frac{1}{3},\frac{2}{3}]$, $\frac{1}{3},\frac{2}{3},-\frac{1}{3}]$, $\frac{2}{3},-\frac{1}{3},-\frac{2}{3}]$ which forms an (ordinary plane) equilateral triangle.
Our operator $\mathcal{T}_{MP}$ (Definition˜1) has an analogous structure, where it is a (scaled) projection operator ($a_{2}P_{\hat{\mathcal{H}}}$) except along the constant function. The $CKA$ (Eq.˜7) also measures the alignment except along the mean direction (for matrices) or the constant direction (for operators). This allows neural networks to have any arbitrary eigenvalue for the constant function, yet perfectly align with $\mathcal{D}_{LR}=0$.
When $\mathcal{T}$ is $\mathcal{T}_{MP}$, the definition of $\mathcal{T}_{MP}$ (Definition˜1) trivially shows that the largest eigenfunction is the constant function. A peculiar observation is that the first eigenfunction of $\mathcal{T}$ remains the constant function ($\braket{\mathbf{1}|e_{1}}>0.95$) throughout training in all our experiments (except Fig.˜9). This pattern also appears in hierarchical datasets (Saxe et al., 2022), suggesting that natural data may consistently prioritize the constant function as the dominant mode — warranting further investigation.